OpenStudy (anonymous):

16. What is the percentage of nitrogen in N2O? Show all calculations leading to your answer.

OpenStudy (photon336):

@JayTheGreat let's do this step by step

OpenStudy (photon336):

1. The first thing we need to do is we must calculate the total mass of our sample.

OpenStudy (anonymous):

ok im writing this down as we go so i can have it just in case i have a similar problem.

OpenStudy (photon336):

2. we do this by counting up the number and types of each atom. 

OpenStudy (photon336):



OpenStudy (photon336):

they asked for the % of nitrogen so we know \[\frac{ mass of nitrogen }{ total mass of compound } = \frac{ N_{2} }{ N_{2}O } = \frac{ 28 }{ 44 } x 100 = 64% N_{2}\]

OpenStudy (photon336):

so @JayTheGreat our compound is 64% nitrogen and 36% oxygen the most important step is the last one, but you'll see that all of the steps are important.

OpenStudy (anonymous):

thank you soooo much. you were a great help, i gave you a medal for your hard work.

OpenStudy (photon336):

no problem @JayTheGreat anytime this is an easy problem but if you like using the steps I showed you tell me the % of hydrogen in water. \[H_{2}O\] if you got it then that's great